OpenStudy (anonymous):

Question 2 Pam likes to practice dancing while preparing for a math tournament. She spends 80 minutes every day practicing dance and math. To help her concentrate better on math, she spends 20 minutes more dancing than doing math. Part A: Write a pair of linear equations to show the relationship between the number of minutes Pam practices math every day (x) and the number of minutes she practices dances every day (y). Part B: How much time does Pam spend on practicing math every day? (3 points) Part C: Is it possible for Pam to have spent 60 minutes practicing dance every day? Explain your reasoning. (2 points)

OpenStudy (anonymous):

nvmd... i found it

OpenStudy (anonymous):

A) m+d=80 d=m+20 B) 30 minutes C) no If Pam was to spend 60 minutes practicing dance, then she would only get 20 minutes of math. 60-20=40. She would be practicing dance 40 minutes more than math, not just 20.